# Establishing Reality-Virtuality Interconnections in Urban Digital Twins for Superior Intelligent Road Inspection

### v1

@@ Abstract
Road inspection is essential for ensuring road maintenance and traffic safety, as road defects gradually emerge and compromise road functionality. Traditional methods, which rely on manual evaluations, are labor-intensive, costly, and time-consuming. Although data-driven approaches are gaining traction, the scarcity and spatial sparsity of road defects in the real world pose significant challenges in acquiring high-quality datasets. Existing simulators designed to generate detailed synthetic driving scenes, however, lack models for road defects. Furthermore, advanced driving tasks involving interactions with road surfaces, such as planning and control in defective areas, remain underexplored. To address these limitations, we propose a system based on Urban Digital Twin (UDT) technology for intelligent road inspection. First, hierarchical road models are constructed from real-world driving data, creating highly detailed representations of road defect structures and surface elevations. Next, digital road twins are generated to create simulation environments for comprehensive analysis and evaluation. These scenarios are subsequently imported into a simulator to enable both data acquisition and physical simulation. Experimental results demonstrate that driving tasks, including perception and decision-making, can be significantly improved using the high-fidelity road defect scenes generated by our system.
Index Terms:
intelligent road inspection, urban digital twins, data acquisition, physical simulationI Introduction
The assessment of road surfaces is critical for vehicle dynamics and driving performance [1]. Road defects, such as cracks and potholes, induce vertical vehicle vibrations, negatively impacting driving comfort and accelerating wear on vehicle components. Timely detection, repair, and maintenance of these defects are essential to ensure road functionality and traffic safety. Current road inspection methods face several challenges [2]. The traditional manual visual inspection process, conducted by certified inspectors, is labor-intensive, time-consuming, and hazardous. Moreover, it significantly disrupts normal traffic flow, making it impractical for large-scale road condition assessments [3, 4, 5]. Although crowdsourcing approaches have been proposed to collect large volumes of defect detection data [6], most of this data consists solely of images, lacking additional sensor modalities. Furthermore, the quality of the collected data is often compromised by individual subjectivity. Given these limitations, there is a growing need for intelligent road inspection systems that are portable, multi-modal, and reliable.
Recent technological advancements have made intelligent road inspection using Urban Digital Twins (UDT) feasible. A UDT system reconstructs the semantic and geospatial properties of urban entities, such as roads and buildings [7]. These digital replicas are indispensable for various applications, including real-time monitoring and simulation. Furthermore, the interconnection between the physical and virtual worlds enables the automatic generation of well-annotated road data, which can be used to train intelligent road inspection models.
Despite these advancements, existing virtual engines, such as CARLA [8] (based on Unreal Engine 4), typically represent roads as 2D planar surfaces. To effectively model road unevenness for perception and planning tasks, it is necessary to reconstruct the 3D structures of simulation assets using real-world data. Given the varying scale of road entities, this challenge can be divided into two parts: creating hierarchical road models using multi-modal data from different sensors, and generating digital road twins for simulation scenarios.
Once a UDT system is developed, a key question arises: can the synthetic data be used to evaluate downstream driving tasks, including perception and decision-making? For perception, models pre-trained on synthetic data demonstrate the ability to generalize to real-world data. Beyond replicating existing scenes, random combinations of road models create novel scenarios that enhance generalization. For planning, we propose a new research area focused on intelligent vehicle systems navigating defected road regions to improve both safety and comfort. Severe road defects can be treated as obstacles in avoidance tasks, requiring collision detection for the vehicle’s wheels rather than its entire body. In scenarios where traversal through defects is unavoidable, speed control can help mitigate vehicle vibrations.
I-A Contributions
@@
To address the aforementioned challenges, we develop an UDT system that enables Reality-Virtuality Interconnection (RVI), specifically tailored for intelligent road inspection. This system not only reconstructs digital twins of road entities for data acquisition, but also generates simulation scenes of defected road models for evaluating perception and decision making methods. The contributions of this study span five key aspects: equipment, algorithm, simulator, dataset, and benchmark. As illustrated in Fig. 1, we develop a portable, multi-sensor experimental setup, equipped with a Livox Mid-360 LiDAR, two FLIR BFS-U3-31S4C cameras, and a DETA100D4G GNSS RTK module. This setup can be mounted on any vehicle, enabling multi-modal, high-quality road data collection in real-world scenarios. The collected data is processed using a LiDAR-Inertial Odometry (LIO) algorithm [9] to construct and continuously update a large-scale 3D map of the road scene. Additionally, a highly accurate stereo matching algorithm [10] is employed to generate dense 3D point clouds of road defects. To model the interconnection between reality and virtuality for urban entities, we develop two key modules: a hierarchical road model creator and a digital road twin generator. The former module extracts road defect structures and undamaged surface elevations from real-world sensor data, while the latter creates high-fidelity scenes for data collection and physical simulation. Specifically, we integrate varying-scaled road entities into a unified road surface model, ensuring no mesh collisions or face isolation for physical simulation. Mesh boundaries are carefully managed to segment 3D models into distinct simulation asset groups for semantic acquisition by the simulator. These assets replace simplified planar road models in the Unreal Engine 4 platform to simulate realistic road scenes. Given the sporadic occurrence of real-world road defects and the challenges inherent in collecting and labeling large-scale, multi-modal, high-quality road data with defects, our simulator can generate an unlimited amount of synthetic road data with realistic structure under diverse weather and lighting conditions. For perception tasks, we create a comprehensive dataset containing semantic and instance-level annotations, along with ground truth data for depth, event, and surface normals, to benchmark State-of-The-Art (SoTA) Deep Neural Networks (DNNs) for intelligent road inspection. Furthermore, unlike current autonomous driving tasks that primarily focus on obstacles impacting the vehicle’s body, our system evaluates decision-making strategies for “negative obstacles”, which consider only wheel-level collisions as constraints. Path planning algorithms are assessed for avoidance, and speed control strategies are analyzed for traversing road defects while maintaining driving comfort. Experimental results demonstrate that the optimal traversing speed varies depending on the specific characteristics of each defect.
@@
We design a UDT system for road inspection, including a portable multi-modal sensor equipment for real-world data collection, and a CARLA-based road defect simulator for data acquisition and physical simulation.
We propose a pipeline to autonomously construct hierarchical road models and generate digital road twins, enabling simulation for driving tasks such as perception, planning, and control in defected road conditions.
We establish a new research paradigm for road scene reconstruction. In addition to generating diverse and well-annotated synthetic data, our system supports driving tasks that physically interact with road surfaces.
The remainder of this article is structured as follows: Sect. II reviews relevant research works. Sect. III and Sect. IV introduce the proposed sensor equipment and UDT system for intelligent road inspection, respectively. Sect. V presents the simulation setup and experimental results for driving tasks using the UDT system. Finally, Sect. VI concludes the study and discusses potential directions for future research.
@@ II-A Intelligent Road Inspection
Traditional road inspection is typically carried out by structural engineers or certified inspectors, which is often subjective, inefficient, expensive, and sometimes hazardous [2]. To address these challenges, researchers have focused on developing intelligent road inspection systems that can recognize, localize, and reconstruct road defects efficiently, accurately, and objectively [11]. Data acquisition is the initial step in intelligent road inspection. Some methods mount sensors on the rear of the vehicle with a downward-facing view, targeting the road surface with a known distance, including single cameras [12, 13], stereo camera pairs [14], laser scanners [15] and laser pointers [16]. However, these specialized sensor configurations are restricted to dedicated inspection vehicles, limiting their scalability, timeliness, and widespread applicability. Alternatively, crowdsourcing approaches can gather extensive data from users’ daily driving. In such setups, sensors are forward-facing, enabling dual functionality for driving perception and road inspection. For instance, [17] employs LiDAR on the hood to detect road unevenness, while [18] and [19] reconstruct forward 3D road surfaces using stereo cameras. Additionally, smartphone applications have emerged as a practical data source for road inspection [20].
Numerous computer vision-based algorithms have been developed for road defect detection. Classical 2D image processing methods have been widely explored, extracting damaged road areas from segmented foregrounds based on geometric and textural assumptions [21, 19]. For instance, to detect obstacles on non-flat road geometries, V-disparity [22] is employed. Subsequently, [10] and [3] refined these representations using disparity transformation algorithms to improve detection accuracy. However, such methods are sensitive to environmental factors that violate their underlying assumptions. 3D point cloud modeling and segmentation have also been employed to detect road irregularities. Camera-based approaches reconstruct dense 3D road point clouds and interpolate them into planar or quadratic surfaces [23, 24, 25]. LiDAR-based methods have been utilized for road roughness perception [17, 26], while road patches traversed by a vehicle’s tires are segmented to identify irregularities such as bumps and potholes [27]. With the rise of machine and deep learning, data-driven approaches have become the dominant techniques. Backbones such as Inception-v2 [28], MobileNet [29], ResNet [30], and RetinaNet [31] have been widely adopted for image classification and object detection for road defects. Image segmentation methods have also been proposed based on the networks, ranging from fully convolutional networks to graph attention-based models [32, 33, 4]. However, the performance of these data-driven methods heavily depends on the quality of the dataset. Due to the rarity and sparse distribution of road defects, capturing sufficient annotated data for training remains a significant challenge. Developing cost-effective methods to address this limitation is critical for advancing intelligent road inspection systems.
@@ II-B Urban Digital Twins
Digital Twin (DT), originally developed for cyber-physical integration in factories, connects real and virtual products through advanced data communication technologies [34, 35]. With the rapid advancement of digital transformation, the applications of DT technology have expanded across various domains, including smart manufacturing [36], advanced medical healthcare[37], and intelligent urban planning [38]. Leveraging advantages such as real-time monitoring [39], fast simulation, [40] and troubleshooting [41], DT technology applied to urban entities integrates data from multiple sensors and creates virtual models to address road anomalies [42]. These capabilities are exemplified by applications such as intelligent road inspection [43]. Extensive research has been conducted on urban entities using digital twin technology. For example, [44] proposed a pavement crack segmentation approach based on 3D edge detection within digital twins. [45] developed a cognitive twin of a pavement structure using data acquired from unmanned aerial vehicles for damage detection. Furthermore, [46] introduced a digital twin model capable of generating virtual distress data for asphalt pavements, while [47] presented a bridge management system leveraging digital twins for precise bridge inspection. Recent studies have also achieved large-scale reconstruction of explicit road surface meshes, recovering both geometry and texture [48, 49]. However, most research treats digital twins of road defects and surfaces as individual reconstruction tasks. In contrast, our work integrates defective road entities at the scene level, creating comprehensive simulation environments specifically tailored for autonomous driving applications.
III Multi-modal Data Acquisition
We employ a portable, multi-sensor, frame-aligned experimental setup (see Fig. 1) to capture various modalities of data. In the setup, the LiDAR generates sparse 3D point clouds, the IMU measures linear acceleration, angular velocity, as well as magnetic field strength, and the stereo camera produces dense disparity maps, where each disparity value is inversely proportional to depth. Rapid sensor movement can cause frame misalignment. Therefore, a GNSS RTK module is used for satellite time correction. A synchronized system clock signal, along with timestamp messages, is broadcast to all sensors to ensure accurate temporal alignment. The pose of the experimental setup is estimated using FAST-LIO [9], a state-of-the-art real-time LiDAR-inertial odometry algorithm. The following prediction model and observation model are utilized for state estimation:
{align}
[…]
where represents the IMU measurements at timestamp , and respectively denote system and observation noises, the system state at timestamp includes position, velocity, rotation, accelerometer bias, gyroscope bias, and the gravity vector, and the observation at timestamp denotes the distance between LiDAR feature points and their corresponding local planes in the estimated map. The FAST-LIO algorithm uses a tightly-coupled iterated extended Kalman filter to update the gain matrix , resulting in the refined error state and predicted state as follows:
{align}
[…]
With pre-estimated extrinsic matrices, sensor frames are spatially and temporally aligned, enabling the acquisition of multi-modal data for our UDT system.
IV Urban Digital Twin System
IV-A System Overview
Our proposed UDT system consists of two main components: a hierarchical road model creator (detailed in Sect. IV-B) and digital road twin generator (detailed in Sect. IV-C).
For the first component, the geometric and texture information of physical entities are extracted from frame-aligned stereo images and LiDAR point clouds, captured using the aforementioned equipment. The hierarchical road mesh creator utilizes this data to generate virtual road models, defined as , where denotes the point clouds within a region of interest captured by the LiDAR and stereo camera, and represents the semantic masks produced using Grounded-SAM [50] and an unsupervised transformed disparity map segmentation algorithm [3]. In existing simulators, such as CARLA, roads are typically modeled as planar surfaces, resulting in meshes that lack sufficient granularity. To accurately represent the unevenness of physical roads, a collection of non-planar meshes must be reconstructed, as illustrated in Fig. 2. Each mesh consists of a set of triangular faces, with each face defined by three vertices and three edges . Generating meshes for an entire road using the stereo camera is computationally complex and memory-intensive, while using LiDAR for this task is impractical for road defects due to the limited granularity of sparse point clouds. When road defects are absent, coarse-grained meshes are preferred, with each vertex corresponding to a 3D point from the sparse point cloud generated by the LiDAR. In contrast, when road defects are present, fine-grained meshes are required, where each vertex corresponds to a 3D point from the dense point cloud generated by the stereo camera. Therefore, we design a hierarchical road model creator with two processing streams, one for road defects and the other for non-defective surfaces.
Subsequently, the digital road twin generator converts the models into simulation assets for scene generation. In the simulator, RGB images are rendered from objects through a pipeline that processes pixel data in stages, such as occlusion culling, vertex shading, and rasterization. To ensure geometric consistency, adjacent meshes must align seamlessly in the scene, with no gaps or overlaps. However, the meshes cannot be modeled as a whole, as semantic annotations require each simulation instance to be labeled with a class ID corresponding to its asset. To provide this reference, the outputs of the hierarchical road model generator are disassembled and saved as individual entities. Finally, a diverse virtual environment is generated using the road entities. Existing simulation scenes, although based on over-simplified planar road surfaces, incorporate built-in navigation features such as traffic lights and reference trajectories. To fully leverage these capabilities, our developed digital road twin generator integrates the detailed defect models with the planar road models and restores the elevation variations of physical roads.
With these two components, our UDT system produces well-annotated road scene data and enhances the comprehensive understanding of road scenes.
IV-B Hierarchical Road Model Creator
@@
The hierarchical road model creator consists of two reconstruction streams, one for road defects and the other for non-defective surfaces, as illustrated in Fig. 3. The coarse-grained stream, focused on the reconstruction of non-defective road surfaces, follows a straightforward process. Semantic masks are generated from camera images using Grounded-SAM, with prompts such as “road” or “pavement”. LiDAR points are segmented based on these masks and fused across multiple frames. To achieve coordinate alignment, a horizontal ground plane is fitted to points from multiple road segments, assuming local elevation variations while maintaining overall flatness. A 3D downsampling operation is then applied to refine the vertices before meshing, accounting for density variations in the point clouds caused by redundant or missing observation views. Given the simple geometry of non-defective road surfaces, elevation can be decoupled from the planar coordinates. The mesh model is reconstructed by directly triangulating the 2D projection of vertices onto the horizontal plane. Vertex elevations are subsequently restored from the original point cloud, ensuring that the reconstructed surface accurately reflects real-world topography.
@@
IV-C Digital Road Twin Generator
Although road scenes can be accurately replicated, road defects are relatively rare in the physical world. To generate more diverse digital road twins, over-simplified road models from the simulator must be integrated with reconstructed road defect models, as illustrated in Fig. 5.
First, the original road segment assets from the simulator, consisting primarily of planar meshes with simple polygonal shapes, are exported. The generated road defect models are then sampled with random poses and scales for integration. When merging the mesh models, intersecting triangular faces must be removed. However, because road segments are not always convex, simply removing all faces and remeshing could compromise their structural integrity.
Therefore, defect models are projected vertically onto the road surface to accurately identify and remove only the intersecting triangles.
The Möller-Trumbore algorithm [51] is typically used to detect ray-triangle intersections,
{align}
[…]
s.t.{0 ≤u ≤1
0 ≤v ≤1
0 ≤1-u-v ≤1
where ray originates from in direction with length , and the triangle is determined by points , , . Rewriting (IV-C) leads to:
{align}
[…]

### v2

@@ Abstract
Road inspection is crucial for maintaining road’s serviceability and ensuring traffic safety, as road defects gradually develop and compromise functionality. Traditional inspection methods, which rely on manual evaluations, are labor-intensive, costly, and time-consuming. While data-driven approaches are gaining traction, the scarcity and spatial sparsity of real-world road defects present significant challenges in acquiring high-quality datasets. Existing simulators designed to generate detailed synthetic driving scenes, however, lack models for road defects. Moreover, advanced driving tasks that involve interactions with road surfaces, such as planning and control in defective areas, remain underexplored. To address these limitations, we propose a multi-modal sensor platform integrated with an urban digital twin (UDT) system for intelligent road inspection. First, hierarchical road models are constructed from real-world driving data collected using vehicle-mounted sensors, resulting in highly detailed representations of road defect structures and surface elevations. Next, digital road twins are generated to create simulation environments for comprehensive analysis and evaluation of algorithm’s performance. These scenarios are then imported into a simulator to facilitate both data acquisition and physical simulation. Experimental results demonstrate that driving tasks, including perception and decision-making, benefit significantly from the high-fidelity road defect scenes generated by our system.
I Introduction
Road condition assessment is essential for ensuring optimal vehicle dynamics and driving performance [guo2018review]. Defects such as cracks and potholes not only induce vibrations but also accelerate the wear of vehicle components [fan2021rethinking]. Timely detection and repair of these defects are therefore crucial for ensuring traffic safety [fan2019road]. Nevertheless, current road inspection methods face significant challenges [kim2014review]. Traditional manual visual inspection, carried out by certified inspectors, is both labor-intensive and hazardous [fan2019pothole]. It also causes substantial disruption to traffic flow, making it impractical for large-scale road condition assessment [wang2024two]. In recent years, data-driven approaches have gained prominence. By fully leveraging extensive driving datasets, deep neural networks (DNNs) can achieve outstanding performance [zhao2023drmnet]. However, despite the unlimited data that can be collected, their effectiveness remains limited by the availability of high-quality ground-truth annotations [yin2024promoting].
Recent technological advancements have enabled the development of intelligent road inspection systems supported by urban digital twins (UDT). A UDT system reconstructs the semantic and geospatial properties of urban entities, such as roads and buildings [guo2024udtiri], providing indispensable digital replicas for real-time monitoring and simulation. However, existing simulators typically represent roads as 2D planar surfaces, unable to capture surface unevenness. As road surfaces directly interact with vehicles and significantly influence driving safety, it is necessary to reconstruct their 3D structures using real-world measurements.
To address these challenges, we develop a UDT system that enables reality-virtuality interconnection (RVI), specifically tailored for intelligent road inspection. The system not only creates digital twins of road entities, but also builds simulation environments containing road defects to support the evaluation of perception and decision-making algorithms. For perception tasks such as semantic scene parsing and 3D geometry reconstruction, it not only replicates real-world road environments but also generates novel scenarios by randomly combining road models to enhance generalization. Experiments show that semantic segmentation and stereo matching networks pre-trained on our synthetic data transfer effectively to real-world datasets. For decision-making tasks, we explore a new paradigm for handling road defects. Instead of following conventional methods that treat defects as obstacles to be avoided entirely, our approach enables either safely bypassing or gliding over certain defects without full detours, leveraging detailed road surface geometry from the UDT. By incorporating tire-level collision detection rather than vehicle-body bounding box checks, the system supports more flexible and efficient obstacle avoidance strategies.
@@ I Introduction
The contributions of this study span five key aspects: equipment, algorithm, simulator, dataset, and benchmark. As illustrated in Fig. 1, we develop a portable, multi-sensor experimental setup, equipped with a Livox Mid-360 LiDAR, two FLIR BFS-U3-31S4C cameras, and a DETA100D4G GNSS RTK module. This setup can be mounted on any vehicle, enabling multi-modal, high-quality road data collection in real-world scenarios. To model the interconnection between reality and virtuality for urban entities, we develop two key modules: a hierarchical road model creator and a digital road twin generator. The former module extracts road defect structures and undamaged surface elevations from real-world sensor data. Real-world road entities across different scales, from coarse-grained road surfaces that shape the overall driving environment to fine-grained defects that directly affect vehicle stability and comfort, are segmented and integrated into unified road surface mesh models, replacing the simplified planar road assets for more realistic road scenes. The latter module creates high-fidelity scenes for synthetic data collection and physical simulation. Given the challenges inherent in large-scale road surface and sporadically occurred road defects in the real world, our system enables seamless integration of any defect model with any surface model smoothly, even if the models originate from different locations. For perception tasks, we create a comprehensive dataset containing semantic and instance-level annotations, along with ground truth data for depth, event, and surface normals, to benchmark state-of-the-art (SoTA) DNNs for intelligent road inspection. For decision-making tasks, we design an experimental framework that treats road defects as negative obstacles, enabling the evaluation of diverse trajectory planning and speed control algorithms. The framework provides physical simulation for both defect-avoidance and traversal strategies that optimize driving comfort, incorporating wheel-level collision constraints for more precise decisions.
@@
We design a UDT system for road inspection, including a portable multi-modal sensor equipment for real-world data collection, and simulation environments containing road defects for data synthesis and physical simulation.
We propose a pipeline to autonomously construct hierarchical road models and generate digital road twins, which both replicate high-fidelity real-world environments and enable the creation of novel simulation scenarios.
We provide a comprehensive benchmark to evaluate perception and decision-making algorithms leveraging our urban digital twins, demonstrating their effectiveness for downstream driving tasks.
The remainder of this article is structured as follows: Sect. II reviews relevant research works. Sect. III introduces the UDT system for intelligent road inspection. Sect. IV presents the simulation setup and experimental results. Finally, Sect. V discusses the limitations and potential directions for future research, while Sect. VI concludes the study.
@@ II-A Intelligent Road Inspection
Traditional road inspection is typically conducted by structural engineers or certified inspectors, a process that is often subjective, inefficient, and sometimes hazardous [kim2014review]. To address these challenges, researchers have developed intelligent road inspection systems capable of automatically collecting data and identifying road defects [ma2022computer]. Among the most direct effects of road defects is vibration. Measurement systems analyze vehicle dynamics as the vehicle traverses road defects [dozsa2022road]. While these systems provide high precision, they are limited to detecting defects within the tire tracks. Other approaches mount visual sensors on the rear of the vehicle with a downward-facing view, maintaining a known distance from the road surface, including single cameras [prasanna2014automated], stereo camera pairs [karunasekera2019energy], and laser scanners [han2012enhanced]. However, such specialized sensor configurations are restricted to dedicated inspection vehicles, limiting scalability and widespread applicability. In contrast, crowd-sourcing approaches collect large-scale data from daily driving using forward-facing sensors, enabling dual functionality for both driving perception and road inspection. For instance, LiDAR-based methods collect data for road unevenness [zhao2023hierarchical], while stereo camera-based approaches reconstruct 3D road surfaces [oniga2009processing]. Smartphone applications have also emerged as a practical data source.
Numerous algorithms have been developed for road defect detection. Classical 2D image processing methods have been widely explored, extracting damaged road areas from segmented foregrounds based on geometric and textural assumptions [fan2021rethinking]. However, such methods are sensitive to environmental factors that violate the underlying assumptions. 3D point cloud methods have also been employed to detect road irregularities. Stereo vision-based approaches reconstruct dense 3D road point clouds and interpolate them into planar or quadratic surfaces. LiDAR-based methods are utilized for road roughness perception [zhao2024roadbev], while road patches traversed by a vehicle’s tires are segmented to identify irregularities such as bumps and potholes [RSRD]. With the rise of deep learning, data-driven approaches have become the dominant techniques [feng2022mafnet]. However, the performance heavily depends on the dataset quality. Due to the rarity and sparse distribution of road defects, capturing sufficient annotated data for training remains a significant challenge.
@@ II-B Urban Digital Twins
Digital twin, originally developed for cyber-physical integration in factories, connects real and virtual products through advanced data communication technologies [lu2020communication]. Leveraging advantages such as real-time monitoring, fast simulation and troubleshooting, digital twin technology applied to urban entities integrates online sensor data and creates virtual models to address road anomalies [wang2024applications]. These capabilities are exemplified by applications such as intelligent road inspection [fan2022urban]. For example, [cao2022pavement] proposed a pavement crack segmentation approach based on 3D edge detection within digital twins. Recent studies have also achieved large-scale reconstruction of explicit road surface meshes, recovering both geometry and texture [mei2024rome, wu2024emie]. However, most research treats digital twins of road defects and surfaces as individual reconstruction tasks. In contrast, our work integrates defective road entities at the scene level, creating comprehensive simulation environments specifically tailored for autonomous driving applications.
III Urban Digital Twin System
III-A Hierarchical Road Model Creator
@@
The hierarchical road model creator consists of two reconstruction streams, one for road defects and the other for non-defective surfaces, as illustrated in Fig. 3. The coarse-grained stream, focused on the reconstruction of non-defective road surfaces, follows a straightforward process. Semantic masks are generated from camera images using Grounded-SAM [groundedsam], with prompts such as “road” or “pavement”. LiDAR points are segmented based on these masks and fused across multiple frames. To achieve coordinate alignment, a horizontal ground plane is fitted to points from multiple road segments, assuming local elevation variations while maintaining overall flatness. A 3D downsampling operation is then applied to refine the vertices before meshing, accounting for density variations in the point clouds caused by redundant or missing observation views. Given the simple geometry of non-defective road surfaces, elevation can be decoupled from the planar coordinates. The mesh model is reconstructed by directly triangulating the 2D projection of vertices onto the horizontal plane. Vertex elevations are subsequently restored from the original point cloud, ensuring that the reconstructed surface accurately reflects real-world topography.
@@
III-B Digital Road Twin Generator
Although road scenes can be accurately replicated, road defects are relatively rare in the physical world. To generate more diverse scenes, over-simplified road models from simulators must be updated with our road defect models, as illustrated in Fig. 5. First, the original planar road segment assets are exported from the simulator to maintain road topology for the scene. The generated road defect models are then sampled with random poses and scales for integration. When merging the mesh models, intersecting triangular faces must be removed. However, because road segments are not always convex, simply removing all faces and re-meshing could compromise their structural integrity. Therefore, defect models are projected vertically onto the road surface to accurately identify and remove only the intersecting triangles. The Möller-Trumbore algorithm [MollerTrumbore], typically used to detect such ray-triangle intersections, is expressed as follows:
where ray originates from in direction with length , and the triangle is determined by points , , and . Rewriting (1) leads to the following expression: